# Optimal time estimation and the clock uncertainty relation for stochastic processes

### v1

@@ I Introduction
While we sincerely hope that our readers never find themselves in this situation, the task of timekeeping using imperfectly periodic events is central to current research aiming to uncover the physical limits on nanoscale clocks in quantum [1, 2, 3, 4, 5, 6, 7], electronic [8, 9, 10, 11], and biomolecular [12, 13, 14] settings. Quantifying the precision of a ticking clock can be seen as an instance of the more general problem of counting statistics in non-equilibrium systems [15]. A major breakthrough in this field was the discovery of the thermodynamic (TUR) [16, 17] and kinetic (KUR) [18, 19] uncertainty relations, which impose universal bounds on fluctuations far from equilibrium. These relations not only limit the accuracy of clocks but also constrain the fluctuations of generic currents and other observables in systems undergoing Markovian stochastic dynamics. Important applications include performance bounds for nanoscale heat engines [20, 21, 22] and methods to infer entropy production from experimentally accessible data [23, 24, 25], while generalisations encompass transient fluctuations [26, 27], first-passage times [18, 28, 29], driven systems [30, 31], ballistic transport [32], quantum dynamics [33, 34, 35, 36, 37], feedback control [38], and relations to fluctuation theorems [39, 40, 41]; see Refs. [42, 43] for reviews.
@@
As explained below, Ineq. (1) emerges naturally from the problem of time estimation, so we term it the clock uncertainty relation (CUR). Like the KUR, the CUR applies to general counting observables, it holds arbitrarily far from equilibrium, and it does not require local detailed balance. Physically, both the KUR and the CUR state that the signal-to-noise ratio is limited by the activity of the process, as measured by $\mathcal{A}$ and $\mathcal{T}^{-1}$, respectively. Unlike the KUR, however, Ineq. (1) can always be saturated and we explicitly construct the counting observables that do so. This distinguishes the inverse mean residual time, $\mathcal{T}^{-1}$, as the proper measure of activity controlling fluctuations in non-equilibrium steady states.
@@
We first prove that this Fisher information, $\mathcal{F}_{t}$, scales inversely to the mean residual time, $\mathcal{F}_{t}\sim 1/(\mathcal{T}t)$, when the time $t$ is large (Sec. II). In our setting, this result yields the ultimate upper bound on time precision when all events and possible correlations between them are accounted for. We also construct an unbiased time estimator that asymptotically saturates the CRB, which is far simpler than the maximum-likelihood estimator because it only involves counting of events. Our presentation is mostly self-contained but some derivations are relegated to the appendices, including details on the formalism of jump trajectories (Appendix A) and current fluctuations (Appendix B) on which many of our other results rely.
We then formulate the CUR (1) for more general observables in non-equilibrium steady states (Sec. III). We prove that this bound can be saturated by an explicit class of counting observables: those that correspond to efficient time estimators. Quantitative comparisons of different precision bounds are presented for random networks and unicyclic clock models, elucidating the conditions under which the TUR, KUR, and CUR differ. We also prove that the KUR is not generally saturable on basic information-theoretic grounds, and explain why the CRB for time estimation gives the tightest bound on fluctuations of counting observables.
@@ II.1 General setup
$P(\bm{\sigma},\mathbf{t}|t)=\frac{W(\tau_{N}|\sigma_{N})}{\Gamma_{\sigma_{N}}}\prod_{j=0}^{N-1}\pi(\sigma_{j+1}|\sigma_{j})W(\tau_{j}|\sigma_{j})p^{{\rm ss}}_{\sigma_{0}},$ (4)
@@ II.2 Residual time versus waiting time and the jump steady state
also known as the forward recurrence time [57]. To understand its meaning, note that if the system is found in state $\sigma$, the mean waiting time until the next jump is $\Gamma_{\sigma}^{-1}=\int_{0}^{\infty}{\rm d}\tau\,W(\tau|\sigma)\tau$  [^1] Technically, this assumes that the total time $t$ is large enough to avoid boundary effects due to the constraint $\sum_{j=0}^{N}\tau_{j}=t$.. (The waiting time preceding any jump $\sigma\to\mu$ is statistically independent of $\mu$; see Appendix A for discussion of this point.) The mean residual time (9) is thus $\mathcal{T}={\rm E}[\tau_{0}]$, i.e. the expected waiting time until the first jump after observation begins at $t_{0}=0$, where the initial state $\sigma$ is sampled from the steady-state distribution, $p_{\sigma}^{\rm ss}$ (see Fig. 2).
@@
To show this, we need the probability that any given jump originates from state $\sigma$. Given that a jump is observed, the probability that the specific transition $\sigma\to\mu$ has occurred is $R_{\mu\sigma}p_{\sigma}^{\rm ss}/\mathcal{A}$ (see Appendix C). Marginalising this over the final state $\mu$ yields the so-called jump steady state [59]
@@
Note that $\mathcal{T}={\rm E}[\Gamma^{-1}_{\sigma}]$ is the arithmetic mean of the conditional waiting times $\{\Gamma^{-1}_{\sigma}\}$ with respect to the steady-state distribution $\mathbf{p}^{\rm ss}$, while $\mathcal{A}^{-1}={\rm E}[\Gamma_{\sigma}]^{-1}$ is the harmonic mean. We therefore have $\mathcal{T}\geq\mathcal{A}^{-1}$, with equality if and only if all escape rates $\Gamma_{\sigma}$ are equal  [^2] This can be proved, for example, using the Cauchy-Schwarz inequality $\mathbf{a}^{\sf T}\mathbf{b}\leq\lVert\mathbf{a}\rVert\cdot\lVert\mathbf{b}\rVert$, with $a_{\sigma}=\sqrt{p^{\rm ss}_{\sigma}\Gamma_{\sigma}}$ and $b_{\sigma}=\sqrt{p^{\rm ss}_{\sigma}/\Gamma_{\sigma}}$. The necessary and sufficient condition for equality is $a_{\sigma}=\mathrm{const.}\times b_{\sigma}$, which implies $\Gamma_{\sigma}=\mathrm{const.}$ whenever $p^{\rm ss}_{\sigma}\neq 0$., i.e. the residual time is greater than the waiting time between jumps on average. This surprising result arises because the unconditional waiting-time distribution, $W(\tau)=\sum_{\sigma}p_{\sigma}^{\rm jss}W(\tau|\sigma)$, is hyper-exponential and thus leads to bunching of jumps, as explained in Appendix D and illustrated in Fig. 2. This is an example of the inspection paradox [47, 48]: the system is likely to be encountered in long-lived states where it spends more time, even if those states appear infrequently in a typical sequence $\bm{\sigma}$.
@@ II.3 Counting states is sufficient for time estimation
The collection of these numbers, $\mathbf{n}(\bm{\sigma})=(n_{1},\ldots,n_{d})$, is known as the empirical distribution [61] and is equivalent to the histogram of the states appearing in the sequence $\bm{\sigma}$. Note that here it is convenient to use the unnormalised empirical distribution as defined in Eq. (14), in contrast to the standard definition [62] $n^{\prime}_{\mu}=n_{\mu}/(N+1)$ which obeys $\sum_{\mu}n^{\prime}_{\mu}=1$  [^3] This should be distinguished from the empirical measure considered in large deviation theory [18], which gives the fraction of total time spent in state $\sigma$. The empirical measure tends to the steady-state distribution $p_{\sigma}^{\rm ss}$ when averaged over a long observation time, while the empirical distribution tends to the jump steady state; see Appendix C for an illustration.
@@ II.4 Asymptotic Fisher information and efficient time estimator
In general, evaluating the Fisher information is complicated, but the problem simplifies in the asymptotic limit of large $t$, where many jumps have occurred. In this limit, the integral in Eq. (13) is dominated by the contribution from the saddle point and we can write (see Appendix E)
@@
Equation (15) is of the form $\partial_{t}\ln P(\bm{\sigma}|t)=f(t)\left[\Theta(\mathbf{n})-t\right]$, where $f(t)$ is independent of the random sequence $\bm{\sigma}$. This is a necessary and sufficient condition for the estimator $\Theta(\mathbf{n})$ to be unbiased and efficient [49], where
$\Theta(\mathbf{n})=\sum_{\sigma=1}^{d}\frac{n_{\sigma}}{\Gamma_{\sigma}}.$ (16)
This estimator has a simple physical interpretation: whenever the state $\sigma$ is observed, one increments the estimate by the expected waiting time until the next jump, $\Theta\to\Theta+\Gamma^{-1}_{\sigma}$. See Fig. 3 for an illustration. Despite its simplicity, this estimator achieves the CRB (7) asymptotically, with the corresponding Fisher information $\mathcal{F}_{t}\approx f(t)$ given by
$\mathcal{F}_{t}=\left(\sum_{\mu=1}^{d}\frac{{\rm E}[n_{\sigma}]}{\Gamma_{\sigma}^{2}}\right)^{-1}+\mathcal{O}(t^{-3/2})=\frac{1}{\mathcal{T}t}+\mathcal{O}(t^{-3/2}).$ (17)
Here, we have used ${\rm E}[n_{\sigma}]=(1+\Gamma_{\sigma}t)p_{\sigma}^{\rm ss}$ (see Appendix F). This is consistent with the definition of the jump steady state (11), as it implies ${\rm E}[n_{\sigma}/(N+1)]=p_{\sigma}^{\rm jss}+\mathcal{O}(t^{-1/2}).$
@@
We can also obtain a simple approximation for the scaled Fisher information in the short-time limit as $t\to 0$. Here, only trajectories with $N=0$ and $N=1$ jumps need be taken into account, and we find (see Appendix E)
@@
In fact, we show in Sec. III.3 [Ineq. (42)] that the upper bound $t\mathcal{F}_{t}\leq\mathcal{A}$ holds for all $t\geq 0$, implying that $t\mathcal{F}_{t}$ is maximal at short times. The corresponding short-time CRB is saturated by the unbiased estimator $\Theta(N)=N/\mathcal{A}$ (see Appendix E). However, this estimator is generally not efficient away from the $t\to 0$ limit, i.e. its variance grows more quickly than $\mathcal{T}t$, in general.
By contrast, the estimator defined in Eq. (16) is only unbiased and efficient in the limit of large $t$, where the saddle-point approximation holds. In particular, as $t\to 0$, it has a bias ${\rm E}[\Theta]={\rm E}[\Gamma_{\sigma}^{-1}]=\mathcal{T}$ and variance ${\rm Var}[\Theta]={\rm Var}[\Gamma_{\sigma}^{-1}]$ (here averages are taken with respect to $\mathbf{p}^{\rm ss}$), as shown in Appendix F and illustrated in Fig. 3. The constant bias can be removed by the trivial shift $\Theta(\mathbf{n})\to\Theta(\mathbf{n})-\mathcal{T}$, although the resulting shifted estimator can take negative values when $t$ is small. However, there exist positive, unbiased, and asymptotically efficient estimators with smaller variance in the short-time limit, as we show in Sec. III.1.
@@
To assess the efficiency of the estimator (16) and the validity of our approximations for the Fisher information, we analyse the example of a two-state system ($d=2$), which can be solved exactly (see Appendix G for details). Figure 4 shows the scaled variance, $\text{Var}[\Theta]/t$, of the estimator in Eq. (16) as a function of time for the two-state system (blue circles in Fig. 4). For comparison, we also plot the inverse scaled Fisher information $1/(t\mathcal{F}_{t})$, computed without any approximation, which represents the lower bound on the scaled variance of any unbiased estimator (red diamonds in Fig. 4). We observe that the scaled variance of the estimator is significantly larger than $1/(t\mathcal{F}_{t})$ for short times. However, after a relatively short time — less than the mean residual time — the variance drops to a value close to the CRB. The inverse scaled Fisher information itself is initially close to $\mathcal{A}^{-1}$ (lower dashed line in Fig. 4) as expected from Eq. (19). Eventually, $1/(t\mathcal{F}_{t})$ tends to the mean residual time $\mathcal{T}$ (upper dashed line in Fig. 4), in agreement with Eq. (18) and confirming the validity of the saddle-point approximation at long times.
@@ III Clock uncertainty relation
Since Ineq. (20) constrains the precision of time estimation, we call it the clock uncertainty relation (CUR). Yet the CUR also bounds the precision of any observable $\Theta(\bm{\sigma})$ that depends on the jump sequence $\bm{\sigma}$ but not the jump times $\mathbf{t}$. The bound is non-trivial for time-extensive observables, such that ${\rm E}[\Theta]\sim t$ and ${\rm Var}[\Theta]\sim t$. This includes the important class of counting observables, defined in Sec. III.1, where we explicitly construct a counting observable that saturates the CUR. We then compare the CUR to other precision bounds in Sec. III.2 and explain its relation to the kinetic uncertainty relation in Sec. III.3.

### v2

@@ I Introduction
While we sincerely hope that our readers never find themselves in this situation, the task of timekeeping using imperfectly periodic events is central to current research aiming to uncover the physical limits on nanoscale clocks in quantum [1, 2, 3, 4, 5, 6, 7], electronic [8, 9, 10, 11], and biomolecular [12, 13, 14] settings. Quantifying the precision of a ticking clock can be seen as an instance of the more general problem of counting statistics in non-equilibrium systems [15]. A major breakthrough in this field was the discovery of the thermodynamic (TUR) [16, 17] and kinetic (KUR) [18, 19] uncertainty relations, which impose universal bounds on fluctuations far from equilibrium. These relations not only limit the accuracy of clocks but also constrain the fluctuations of generic currents and other observables in systems undergoing Markovian stochastic dynamics. Important applications include performance bounds for nanoscale heat engines [20, 21, 22] and methods to infer entropy production from experimentally accessible data [23, 24, 25], while generalisations encompass transient fluctuations [26, 27], first-passage times [18, 28, 29], driven systems [30, 31], ballistic transport [32], quantum dynamics [33, 34, 35, 36, 37, 38], feedback control [39], general counting observables [40], and relations to fluctuation theorems [41, 42, 43]; see Refs. [44, 45] for reviews.
@@
As explained below, Ineq. (1) emerges naturally from the problem of time estimation, so we term it the clock uncertainty relation (CUR). Like the KUR, the CUR applies to general counting observables, it holds arbitrarily far from equilibrium, and it does not require local detailed balance. Physically, both the KUR and the CUR state that the signal-to-noise ratio is limited by the activity of the process, as measured by $\mathcal{A}$ and $\mathcal{T}^{-1}$, respectively. Unlike the KUR, however, Ineq. (1) can always be saturated and we explicitly construct the counting observables that do so. This distinguishes the inverse mean residual time, $\mathcal{T}^{-1}$, as the proper measure of activity controlling fluctuations in non-equilibrium steady states. Moreover, since $\mathcal{T}$ is directly observable and the CUR bound is tight, it can be used as a witness to verify or falsify that the observed dynamics is autonomous, classical, and memoryless.
@@
We first prove that this Fisher information, $\mathcal{F}_{t}$, scales inversely to the mean residual time, $\mathcal{F}_{t}\sim 1/(\mathcal{T}t)$, when the time $t$ is large (Sec. II). In our setting, this result yields the ultimate upper bound on time precision when all events and possible correlations between them are accounted for. We also construct an unbiased time estimator that asymptotically saturates the CRB, which is far simpler than the maximum-likelihood estimator because it only involves counting of events.
We then formulate the CUR (1) for more general observables in non-equilibrium steady states (Sec. III). We prove that this bound can be saturated by an explicit class of counting observables: those that correspond to efficient time estimators. Quantitative comparisons of different precision bounds are presented for random networks and unicyclic clock models, elucidating the conditions under which the TUR, KUR, and CUR differ. We also prove that the KUR is not generally saturable on basic information-theoretic grounds, and explain why the CRB for time estimation gives the tightest bound on fluctuations of counting observables. As an application of the CUR, we show that its experimental violation can be used to detect the presence of unobservable states or transitions, a common scenario in many biological settings [55, 56, 57]. This provides a novel tool for characterising stationary stochastic systems, complementing recent proposals to detect hidden states from transient behaviour [53] or infer entropy production from partially observed dynamics [58, 59, 60].
@@ II.1 General setup
$P(\bm{\sigma},\mathbf{t}|t)=\frac{W(\tau_{N}|\sigma_{N})}{\Gamma_{\sigma_{N}}}\left(\prod_{j=0}^{N-1}\pi(\sigma_{j+1}|\sigma_{j})W(\tau_{j}|\sigma_{j})\right)p^{{\rm ss}}_{\sigma_{0}},$ (4)
@@ II.2 Residual time versus waiting time and the jump steady state
also known as the forward recurrence time [65] or the persistence time [66]. To understand its meaning, note that if the system is found in state $\sigma$, the mean waiting time until the next jump is $\Gamma_{\sigma}^{-1}=\int_{0}^{\infty}{\rm d}\tau\,W(\tau|\sigma)\tau$  [^1] Technically, this assumes that the total time $t$ is large enough to avoid boundary effects due to the constraint $\sum_{j=0}^{N}\tau_{j}=t$.. (The waiting time preceding any jump $\sigma\to\mu$ is statistically independent of $\mu$; see Appendix A for discussion of this point.) The mean residual time (9) is thus $\mathcal{T}={\rm E}[\tau_{0}]$, i.e. the expected waiting time until the first jump after observation begins at $t_{0}=0$, where the initial state $\sigma$ is sampled from the steady-state distribution, $p_{\sigma}^{\rm ss}$ (see Fig. 2).
@@
To show this, we need the probability that any given jump originates from state $\sigma$. Given that a jump is observed, the probability that the specific transition $\sigma\to\mu$ has occurred is $R_{\mu\sigma}p_{\sigma}^{\rm ss}/\mathcal{A}$ (see Appendix B). Marginalising this over the final state $\mu$ yields the so-called jump steady state [68]
@@
Note that $\mathcal{T}={\rm E}[\Gamma^{-1}_{\sigma}]$ is the arithmetic mean of the conditional waiting times $\{\Gamma^{-1}_{\sigma}\}$ with respect to the steady-state distribution $\mathbf{p}^{\rm ss}$, while $\mathcal{A}^{-1}={\rm E}[\Gamma_{\sigma}]^{-1}$ is the harmonic mean. We therefore have $\mathcal{T}\geq\mathcal{A}^{-1}$, with equality if and only if all escape rates $\Gamma_{\sigma}$ are equal  [^2] This can be proved, for example, using the Cauchy-Schwarz inequality $\mathbf{a}^{\sf T}\mathbf{b}\leq\lVert\mathbf{a}\rVert\cdot\lVert\mathbf{b}\rVert$, with $a_{\sigma}=\sqrt{p^{\rm ss}_{\sigma}\Gamma_{\sigma}}$ and $b_{\sigma}=\sqrt{p^{\rm ss}_{\sigma}/\Gamma_{\sigma}}$. The necessary and sufficient condition for equality is $a_{\sigma}=\mathrm{const.}\times b_{\sigma}$, which implies $\Gamma_{\sigma}=\mathrm{const.}$ whenever $p^{\rm ss}_{\sigma}\neq 0$., i.e. the residual time is greater than the waiting time between jumps on average. This surprising result arises because the unconditional waiting-time distribution, $W(\tau)=\sum_{\sigma}p_{\sigma}^{\rm jss}W(\tau|\sigma)$, is hyper-exponential and thus leads to bunching of jumps, as explained in Appendix C and illustrated in Fig. 2. This is an example of the inspection paradox [49, 50]: the system is likely to be encountered in long-lived states where it spends more time, even if those states appear infrequently in a typical sequence $\bm{\sigma}$.
@@ II.3 Counting states is sufficient for time estimation
The collection of these numbers, $\mathbf{n}(\bm{\sigma})=(n_{1},\ldots,n_{d})$, is known as the empirical distribution [70] and is equivalent to the histogram of the states appearing in the sequence $\bm{\sigma}$. Note that here it is convenient to use the unnormalised empirical distribution as defined in Eq. (14), in contrast to the standard definition [71] $n^{\prime}_{\mu}=n_{\mu}/(N+1)$ which obeys $\sum_{\mu}n^{\prime}_{\mu}=1$  [^3] This should be distinguished from the empirical measure considered in large deviation theory [18], which gives the fraction of total time spent in state $\sigma$. The empirical measure tends to the steady-state distribution $p_{\sigma}^{\rm ss}$ when averaged over a long observation time, while the empirical distribution tends to the jump steady state; see Appendix B for an illustration.
@@ II.4 Asymptotic Fisher information and efficient time estimator
In general, evaluating the Fisher information is complicated, but the problem simplifies in the asymptotic limit of large $t$, where many jumps have occurred. In this limit, the integral in Eq. (13) is dominated by the contribution from the saddle point and we can write (see Appendix D)
@@
Equation (15) is of the form $\partial_{t}\ln P(\bm{\sigma}|t)=f(t)\left[\Theta(\mathbf{n})-t\right]$, where $f(t)$ is independent of the random sequence $\bm{\sigma}$. This is the necessary and sufficient condition for the estimator
$\Theta(\mathbf{n})=\sum_{\sigma=1}^{d}\frac{n_{\sigma}}{\Gamma_{\sigma}}$ (16)
to be unbiased and efficient, with $f(t)$ equal to the Fisher information [51]. The efficient estimator in Eq. (16) has a simple physical interpretation: whenever the state $\sigma$ is observed, one increments the estimate by the expected waiting time until the next jump, $\Theta\to\Theta+\Gamma^{-1}_{\sigma}$; see Fig. 3 for an illustration. Despite its simplicity, this estimator achieves the CRB (7) asymptotically, with the corresponding Fisher information given by the prefactor of Eq. (15):
$\mathcal{F}_{t}=\left(\sum_{\sigma=1}^{d}\frac{{\rm E}[n_{\sigma}]}{\Gamma_{\sigma}^{2}}\right)^{-1}+\mathcal{O}(t^{-3/2})=\frac{1}{\mathcal{T}t}+\mathcal{O}(t^{-3/2}).$ (17)
Here, we have used ${\rm E}[n_{\sigma}]=(1+\Gamma_{\sigma}t)p_{\sigma}^{\rm ss}$ (see Appendix E). This is consistent with the definition of the jump steady state (11), as it implies ${\rm E}[n_{\sigma}/(N+1)]=p_{\sigma}^{\rm jss}+\mathcal{O}(t^{-1/2}).$
@@
We can also obtain a simple approximation for the scaled Fisher information in the short-time limit as $t\to 0$. Here, only trajectories with $N=0$ and $N=1$ jumps need be taken into account, and we find (see Appendix D)
@@
In fact, we show in Sec. III.3 [Ineq. (42)] that the upper bound $t\mathcal{F}_{t}\leq\mathcal{A}$ holds for all $t\geq 0$, implying that $t\mathcal{F}_{t}$ is maximal at short times. The corresponding short-time CRB is saturated by the unbiased estimator $\Theta(N)=N/\mathcal{A}$ (see Appendix D). However, this estimator is generally not efficient away from the $t\to 0$ limit, i.e. its variance grows more quickly than $\mathcal{T}t$, in general. By contrast, the estimator defined in Eq. (16) is only unbiased and efficient in the limit of large $t$, where the saddle-point approximation holds. In particular, as $t\to 0$, it has a bias ${\rm E}[\Theta]={\rm E}[\Gamma_{\sigma}^{-1}]=\mathcal{T}$ and variance ${\rm Var}[\Theta]={\rm Var}[\Gamma_{\sigma}^{-1}]$ (here averages are taken with respect to $\mathbf{p}^{\rm ss}$), as shown in Appendix E and illustrated in Fig. 3. The constant bias can be removed by the trivial shift $\Theta(\mathbf{n})\to\Theta(\mathbf{n})-\mathcal{T}$, although the resulting shifted estimator can take negative values when $t$ is small. However, there exist positive, unbiased, and asymptotically efficient estimators with smaller variance in the short-time limit, as we show in Sec. III.1.
@@
To assess the efficiency of the estimator (16) and the validity of our approximations for the Fisher information, we analyse the example of a two-state system ($d=2$), which can be solved exactly (see Appendix F for details). Figure 4 shows the scaled variance, $\text{Var}[\Theta]/t$, of the estimator in Eq. (16) as a function of time for the two-state system (blue circles in Fig. 4). For comparison, we also plot the inverse scaled Fisher information $1/(t\mathcal{F}_{t})$, computed without any approximation, which represents the lower bound on the scaled variance of any unbiased estimator (red diamonds in Fig. 4). We observe that the scaled variance of the estimator is significantly larger than $1/(t\mathcal{F}_{t})$ for short times. However, after a relatively short time — less than the mean residual time — the variance drops to a value close to the CRB. The inverse scaled Fisher information itself is initially close to $\mathcal{A}^{-1}$ (lower dashed line in Fig. 4) as expected from Eq. (19). Eventually, $1/(t\mathcal{F}_{t})$ tends to the mean residual time $\mathcal{T}$ (upper dashed line in Fig. 4), in agreement with Eq. (18) and confirming the validity of the saddle-point approximation at long times.
@@ III Clock uncertainty relation
Since Ineq. (20) constrains the precision of time estimation, we call it the clock uncertainty relation (CUR). Yet the CUR also bounds the precision of any observable $\Theta(\bm{\sigma})$ that depends on the jump sequence $\bm{\sigma}$ but not the jump times $\mathbf{t}$. The bound is non-trivial for time-extensive observables, such that ${\rm E}[\Theta]\sim t$ and ${\rm Var}[\Theta]\sim t$. This includes the important class of counting observables, defined in Sec. III.1, where we explicitly construct a counting observable that saturates the CUR. We then compare the CUR to other precision bounds in Sec. III.2 and explain its relation to the kinetic uncertainty relation in Sec. III.3. Sec. III.4 concludes with an application of the CUR: the detection of hidden transitions from partial observations of steady-state jump and waiting-time statistics.